OpenStudy (anonymous):

(3x/4) to the 3rd power

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

\[\LARGE \left( \frac{3x}{4}\right)^3=\frac{27x^3}{64}\]

OpenStudy (anonymous):

this becomes 27x^3/64

OpenStudy (anonymous):

fast bro :P

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

you too ;)

OpenStudy (anonymous):

mutual medals? :]

OpenStudy (anonymous):

why you're all after medals? ... skipping levels doesn't make you clever ! lol